OpenStudy (anonymous):

find the slope of the function's graph at the given point, then find and equation for the line tangent to the graph f(x)=x-2x^2 , (1,-1)

OpenStudy (anonymous):

start by solving for the slope of the equation

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

i have to use this formula: y=mx+b?

OpenStudy (anonymous):

or y2-y1/ x2-x1?

OpenStudy (anonymous):

x= 1?

OpenStudy (anonymous):

ok so now use the formula (y-y1)=m(x-x1) where (x1,y1)=(1,-1)

OpenStudy (anonymous):

slope=1?

OpenStudy (anonymous):

no your slope will be m

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

using that equation i get y= mx-2

OpenStudy (anonymous):

so y=x-2

OpenStudy (anonymous):

okay,

OpenStudy (anonymous):

now we need to find the inverse of the line

OpenStudy (anonymous):

wait are you sure that you wrote the initial equation right because that is a hyperbola

OpenStudy (anonymous):

let me check again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it's y= mx-2

OpenStudy (anonymous):

for the origional equation f(x)=x-2x^2

OpenStudy (anonymous):

yes, its correct

OpenStudy (anonymous):

ok well either way now we need to find the equation of a tangent line

OpenStudy (anonymous):

how do I that?

OpenStudy (anonymous):

1 sec g2g take a quiz b back in 10

OpenStudy (anonymous):

okay,np

OpenStudy (anonymous):

hmm after thinking about it i think i might have done this wrong

OpenStudy (anonymous):

brb